What is the freezing point of a solution of 10 grams of glucose(C6H12O6) and 1500 grams of water


1
Expert's answer
2021-05-31T02:07:41-0400

Molality = 0.056/(1.5)= 0.037

put the values,


0- t2= 1×1.86×0.037 = 0.07


Freezing Point = 0.07°C